A resident individual investor sold the following shares during the year which have all been held more than 12 months and none are exempt:

Share parcel

Cost base

$

Indexed cost base

$

Sale price

$

A

25,000

26,000

75,000

B

30,000

40,000

44,000

What is their minimum possible net capital gain ?

Group of answer choices

$56,000

$29,000

$32,000

$53,000
Answer #1

The correct answer is 53000

Capital Gain In case of A=Sale value-Index cost=75000-26000=49000

Capital gain in case of B=Sale Value-Index cost=44000-4000=4000

Total =53000